OpenStudy (anonymous):

(1,-3);y+2=4(x-1) plz help the answer is y=4x-7 but i got y=4x-8 how

jigglypuff314 (jigglypuff314):

y + 2 = 4(x - 1) first distribute y + 2 = 4x - 4 then subtract 2 from both sides -2 -2 y = 4x - 8

OpenStudy (therealmeeeee):

slope= 2/3 y-intercept= -1 2. slope = -1/1 y-intercept= 4 Mark the y-intercept (the point that intersects with the y-axis) on the graph. Slope = the ratio of rise over run. So on the first one you would go up 2 units from (0, -1) and then right 3 units and that would be another point (3, 1) on the line. Then draw a line through those two points. On the next one you would go down 1 unit from (0, 4) and then right 1 unit and that would be another point (1, 5) on the line. Then draw a line through those two points.

OpenStudy (anonymous):

yes but the answer from the text book is y=4x-7 is the textbook worng?

OpenStudy (therealmeeeee):

y -2/3x = -1 y -x = 4 -------------------- substitute y= 2/3x-1 in the second equation we get 2/3x-1 -x =4 2/3x-x = 5 x(2/3-1)=5 x(-1/3)=5 x= -5 *(3/1) x= -15 substitute the value in y= 2/3x -1 we get y = 2/3 *(-15)-1 =-10 -1 y=-11 the both lines meet at (-15,-11)

OpenStudy (therealmeeeee):

did any of my stuff help u

OpenStudy (anonymous):

it got me confused

OpenStudy (therealmeeeee):

wattt X_X

OpenStudy (therealmeeeee):

wat part

jigglypuff314 (jigglypuff314):

I don't see how they could have gotten y = 4x - 7 ...

OpenStudy (therealmeeeee):

who me

jigglypuff314 (jigglypuff314):

meeeee, you're work is confusing... the question was, why does the text book say y=4x-7 when the person is getting y=4x-8 all I can come up with is that when you plug in the coordinate in the beginning (1,-3) it proves true for the y=4x-7 --> (-3)=4(1)-7 -> -3=-3

OpenStudy (therealmeeeee):

oh